OpenStudy (anonymous):

How would this problem be worked out? 5(2x+1)=12x-1

OpenStudy (anonymous):

Alright, what's the first step?

OpenStudy (anonymous):

5(2x) and 5(1) (10x+5)=12x-1

OpenStudy (anonymous):

Sweet! So we have: 10x + 5 = 12x - 1 What's next?

OpenStudy (anonymous):

- 5 on both sides?

OpenStudy (anonymous):

Yep! So what do you get?

OpenStudy (anonymous):

10x=12x(6)

OpenStudy (anonymous):

Not quite. You're right on the 10x side, but on the other side, don't forget that you're subtracting, not multiplying. So what's it look like?

OpenStudy (anonymous):

10x=12x(-6)

OpenStudy (anonymous):

No parenthesis (it makes multiplication), so it'll be: 10x =12x - 6 So what's the next step?

OpenStudy (anonymous):

oh ok! +6 on both sides

OpenStudy (anonymous):

Nope, let's get the Xs together. So how do we do that?

OpenStudy (anonymous):

divide by 10x on both sides

OpenStudy (anonymous):

Nope, let's put the 12x on the left side. How would you do that?

OpenStudy (anonymous):

x's 12 on both sides?

OpenStudy (anonymous):

You can do it just like before, just with the 12x, so subtract the12x from the right and subtract it from the left, too: 10x - 12x =12x - 12x - 6 So now what does it look like?

OpenStudy (anonymous):

-2x=-6 divide by -2 on both sides left side cancles out x=-8

OpenStudy (anonymous):

Okay so we have this: -2x = -6 Now when you divide by -2 on the left you just get x by itself (like you said.) But what happens when you take -6 and divide it by -2?

OpenStudy (anonymous):

you get -8??

OpenStudy (anonymous):

Remember that you're dividing.

OpenStudy (anonymous):

What's \[\frac{ -6 }{ -2 }\] (and also what happens when you divide a negative number by another negative number?)

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

6/2 = 3 And when you take -6 and divide it by -2, you get positive 3 (because a negative divided by a negative is a positive.) Do you see why?

OpenStudy (anonymous):

yes i see y its a positive its wierd too lol cause i did -6/-2 on the calc. and -8 everytime lol but now i see

OpenStudy (anonymous):

If you add -6 and -2 you get -8. Always be careful with the arithmetic. It's seriously mess it up every time on a test. The idea behind what I'm doing is usually correct, but I mess it up because I add or do something wrong with my arithmetic. Good thing there's partial credit!

OpenStudy (anonymous):

haha yeahi know right! I'm nowhere good at math, im so behind in it like all this im posting is simple stuff and i never really learned cause i goofed off so much throughout highschool and now im having to take it online as a half senior /junior im graduating early math just really holds me back i really regret goofing off in the past lol so thanks alot for making do the work so i can learn how to it right and not just telling me the anwser right away you've been a real bi help(:

OpenStudy (anonymous):

Oh, no worries. It's a pleasure. Feel free to tag me in any questions you have.